# From Complex–Analytic Models to Dyadic Methods: A Real–Variable Approach to Hypersingular Operators

### v5

@@ Abstract.
Motivated by the work of Cheng–Fang–Wang–Yu on the hypersingular Bergman projection, we develop a real-variable, dyadic framework for hypersingular operators in regimes where strong-type bounds fail on the critical line. Our main new input is what we call the Forelli–Rudin method: a dyadic real-variable mechanism, inspired by complex-analytic Forelli–Rudin type arguments, that provides a flexible way to obtain critical-line and endpoint estimates for general hypersingular operators.
In the unit disc setting with $1<t<3/2$, we obtain a full characterization of the $(p,q)$ mapping theory for the dyadic hypersingular maximal operator $\mathcal{M}_{t}^{\mathcal{D}}$, in particular including estimates on the critical line 1/q1/p=2t21/q-1/p=2t-2 and a weighted endpoint criterion in the radial setting. In addition, we establish a novel two-weight estimate for $\mathcal{M}_{t}^{\mathcal{D}}$ in the range $p>q$, valid for any $t>0$. We also prove endpoint estimates for the hypersingular Bergman projection
@@
with particular emphasis on the weak-type bound at the endpoint $(p,q)=\bigl(\tfrac{1}{3-2t},1\bigr)$. Finally, we introduce a class of hypersingular cousin of sparse operators in $\mathbb{R}^{n}$ associated with graded sparse families, quantified by the sparseness $\eta$ and a new structural parameter (the degree) $K_{\mathcal{S}}$, and we characterize the corresponding strong and weak type regimes in terms of $(n,t,\eta,K_{\mathcal{S}})$.
Our real-variable perspective addresses an inquiry raised by Cheng–Fang–Wang–Yu on developing effective real-analytic tools in the hypersingular regime for both ${\mathcal{M}}_{t}^{{\mathcal{D}}}$ and $K_{2t}$, and it also provides a new route toward the critical-line analysis of Forelli–Rudin type operators and related hypersingular operators in both real and complex settings.
@@ 3. $\bullet$
We have the following full characterization for the behavior of ${\mathcal{M}}_{t}^{{\mathcal{D}}}$, including the critical line estimates. Let $1<t<3/2$ and $\mathcal{D}$} be any dyadic system on $\mathbb{T}$. Then ${\mathcal{M}}_{t}^{{\mathcal{D}}}$ (as well as ${\mathcal{M}}_{t}$) is
@@
The results above are summarized in Figure 1 below.
@@
We summarize the above results for $K_{2t}$ in Figure 2 below.
@@
${\mathcal{S}}_{m}:=\left\{I_{k}:=\left[\frac{k}{2^{m}},\frac{k+1}{2^{m}}\right),0\leq k\leq 2^{m}-1\right\}\cup\left\{[0,2)\right\}.$
@@
which implies $\left\|\mathbb{A}_{{\mathcal{S}}_{m}}^{t}\right\|_{L^{\infty}([0.2))\to L^{1}([0,2))}\gtrsim 2^{m(t-1)}$. This example shows that, in order to capture the behavior of $\mathbb{A}_{{\mathcal{S}}}^{t}$, the sparseness of $\mathcal{S}$} alone does not suffice, and one must have certain control for the dyadic scales between consectutive layers in $\mathcal{S}$}. This simple example also highlights a fundamental difference between the sparse operator and its hypersingular counterpart: in the former case, the degree plays essentially no role. We refer the reader to Section 5.1 for further discussion of this topic.
@@
The difficulties at this endpoint are twofold:
@@
    The strong-type range above is sharp, as can be seen from the strong-type characterization for $K_{2t}$ (see [14, Theorem 3]). These observations suggest that the $L^{p}$$L^{q}$ mapping properties of $K_{2t}$ are governed primarily by the geometry of (dyadic) Carleson boxes, rather than analyticity.
@@
    From the viewpoint of complex function theory, we develop a systematic theory of the hypersingular maximal operator ${\mathcal{M}}_{t}^{{\mathcal{D}}}$ and also advance the study of Forelli–Rudin type operators in the hypersingular regime by establishing new endpoint and critical-line estimates for $K_{2t}$ with $t>1$. To the best of our knowledge, these estimates are new even in the unit disc setting. A key feature of our approach is the introduction of Forelli–Rudin method into this setting, which allows us to overcome the lack of strong-type bounds and provides a robust harmonic-analytic framework for hypersingular operators in both real and complex settings.
@@
The rest of this paper is organized as follows. In Section 2, we collect basic definitions and notation that will be used throughout the paper. In Section 3, we establish the $L^{p}$ mapping theory for the dyadic hypersingular maximal operator $M_{t}^{\mathcal{D}}$, including both the off-critical-line bounds and the critical-line estimates, together with the endpoint characterizations in the radial weighted setting. Moreover, we prove a two-weight estimate for ${\mathcal{M}}_{t}^{{\mathcal{D}}},t>0$. In Section 4, we turn to the hypersingular Bergman projection $K_{2t}$ and prove the critical-line estimates, with special emphasis on the endpoint $(p,q)=\left(\frac{1}{3-2t},1\right)$ via a complex–analytic version of the Forelli–Rudin method. In Section 5, we introduce hypersingular sparse operators associated with graded sparse families in $\mathbb{R}^{n}$ and prove the corresponding $L^{p}$-$L^{q}$ theory via a dyadic version of the Forelli–Rudin method, which in particular extends and strengthens the results obtained earlier for $K_{2t}$. In addition, we include a discussion of two other different approaches to study $\mathbb{A}_{{\mathcal{S}}}^{t}$: one via ideas from sparse domination, and the other via Bourgain’s interpolation trick. Finally, in Section 6, we propose several open problems and further directions motivated by the methods developed in this paper.
@@ 2. Notations
$\left\|f\right\|_{L^{p,r}({\mathbb{D}})}:=\left(p\int_{0}^{\infty}t^{r}\left|\left\{z\in{\mathbb{D}}:|f(z)|>\lambda\right\}\right|^{\frac{r}{p}}\frac{d\lambda}{\lambda}\right)^{\frac{1}{r}},\qquad 0<r<+\infty,$
@@ 3.2. Critical line estimate
$2^{-k}$ $\simeq\left|\bigcup_{I\in{\mathcal{D}}_{k}}Q_{I}\right|\lesssim\left|\left\{w\in{\mathbb{D}}:{\mathcal{M}}_{t}^{{\mathcal{D}}}f_{k}(w)>c2^{k(2t-1)}\cdot\int_{1-\frac{1}{2^{k}}}^{1-\frac{1}{2^{k+1}}}\frac{dr}{\omega^{\frac{3-2t}{2t-2}}(r)}\right\}\right|$
@@ 3.3. A two-weight estimate for $\mathcal{M}_{t}^{\mathcal{D}}$
Let $1\leq q<p\leq+\infty$ and $t>0$. Let $\mathcal{D}$ be a dyadic system on $\mathbb{T}$. Let $\mu$ and $\omega$ be two weights on $\mathbb{D}$, and set $\sigma:=\omega^{-1/(p-1)}$. Assume that $\mu,\sigma\in{\mathcal{B}}_{\infty}({\mathcal{D}})$. Then the following statements are equivalent.
@@
$(1)\Longrightarrow(2)$.   Let $\{a_{I}\}_{I\in{\mathcal{D}}}$ be any sequence of positive numbers, and
@@
${\mathcal{M}}_{t}^{{\mathcal{D}}}F(z)\geq\frac{1}{|Q_{I}|^{t}}\int_{Q_{I}}|F(z)|dA(z)\geq\frac{1}{|Q_{I}|^{t}}\int_{Q^{\textrm{up}}_{I}}|F(z)|dA(z)=\frac{a_{I}\sigma(Q_{I}^{\textrm{up}})}{|Q_{j}|^{t}}\gtrsim\frac{a_{I}\sigma(Q_{I})}{|Q_{j}|^{t}},$
@@
$\int_{{\mathbb{D}}}\left|{\mathcal{M}}_{t}^{{\mathcal{D}}}F(z)\right|^{q}\mu(z)dA(z)$ $\gtrsim\sum_{I\in{\mathcal{D}}}\int_{Q_{I}^{\textrm{up}}}\left(\frac{a_{I}\sigma(Q_{I})}{|Q_{j}|^{t}}\right)^{q}\mu(z)dA(z)$
@@ 4. Critical line estimates for the hypersingular Bergman projection
$\left\{\left(\frac{1}{p},\frac{1}{q}\right)\in[0,1]^{2}:\frac{1}{q}-\frac{1}{p}=2t-2\right\}.$
@@
Therefore, it remains to verify claims (1) and (2). Claim (1) follows from [35, Theorem 3] by choosing $a=2t-2,b=0,c=2t,n=1,\alpha=\beta=0$, and $p=q=\frac{1}{3-2t}$. Claim (2) follows from a direct computation. ∎
@@ 5. Hypersingular sparse operators
    $E(Q)\cap Q^{\prime}=\emptyset$ for any $^{\prime}\subset Q, QQ$$^{\prime}\in Q$.
@@ 5.1. Graded family
We now turn to some details. Let ${\mathcal{G}}\subseteq{\mathcal{D}}$ be a collection of dyadic cubes. Again, we may assume $Q\subseteq[0,1]^{n}$ for all $Q\in{\mathcal{S}}$.
@@
We call $K_{{\mathcal{G}}}$ the degree of $\mathcal{G}$}, and we refer to ${\mathcal{G}}_{j}$ as the $j$-th layer of $\mathcal{G}$}.
@@
    Here, we may assume that ${\mathcal{G}}_{1}\neq\emptyset$. Otherwise, $\mathbb{A}_{{\mathcal{G}}}^{t}f(x)=\mbox{1\hskip-4.25pt{1}}_{[0,1]^{n}}(x)\int_{[0,1]^{n}}f$, which is a rank-one operator and, in particular, maps $L^{1}(\mathbb{R}^{n})$ boundedly into $L^{\infty}(\mathbb{R}^{n})$.
@@
    Here are some examples of graded family. Let ${\mathcal{G}}^{(1)}$ consist of $0,1]^{n}$ together with all of its dyadic descendants; then ${\mathcal{G}}^{(1)}$ is graded (in fact, its degree is $1$). Another example is the collection [^5] In that setting, the role of $Q_{0}$ is replaced by $\mathbb{D}$}, and the sidelength $\ell(Q)$ in Definition 5.5 is replaced by the length of the boundary arc associated with the Carleson box. of dyadic Carleson boxes that appeared in our earlier analysis of ${\mathcal{M}}_{t}^{{\mathcal{D}}}$ and $K_{2t}$. This family is also graded, again with degree $1$. Finally, we observe that $K_{{\mathcal{G}}}\geq 1$.
@@
    (Critical line estimate) $\mathbb{A}_{{\mathcal{S}}}^{t}$ extends to a bounded operator from $L^{p,\infty}(\mathbb{R}^{n})$ to $L^{q}(\mathbb{R}^{n})$ when $(p,q)$ belongs to
@@
Theorem 5.7 strengthens our earlier results on the boundedness behavior of $K_{2t}$ (see, Section 4). Moreover, the proof of Step II in Theorem 5.7 yields an alternative proof of [14, Theorem 3], and the proof of Step III in in Theorem 5.7 offers a different proof of Proposition 4.2.
@@
respectively. We have the following claims.
@@
$\mathbb{A}_{{\mathcal{S}}}^{t}f(x)\lesssim{\bf W}(x)M_{\textrm{HL}}^{{\mathcal{D}}}f(x),\qquad x\in[0,1]^{n},$
@@
(5.9) $\mathbb{A}_{{\mathcal{S}}}^{t}f(x)=\sum_{x\in Q\in{\mathcal{S}}}|Q|^{1-t}\cdot\frac{1}{|Q|}\int_{Q}|f(y)|dy\leq M_{\textrm{HL}}^{{\mathcal{D}}}f(x)\cdot\sum_{x\in Q\in{\mathcal{S}}}|Q|^{1-t}.$
@@
(5.10) $\sum_{x\in Q\in{\mathcal{S}}}|Q|^{1-t}\leq\sum_{j=0}^{{\bf N}(x)}2^{njK_{{\mathcal{S}}}(t-1)}\lesssim 2^{n{\bf N}(x)K_{{\mathcal{S}}}(t-1)}={\bf W}(x).$
@@
and let $f\in L^{p}(\mathbb{R}^{n})$. Without loss of generality, we may assume $\textrm{supp}f\subseteq[0,1]^{n}$. By [17, Exercise 1.4.14] and using the assumption that ${\mathbb{A}}_{{\mathcal{S}}}^{t}$ is localized in $0,1]^{n}$, we have
@@
$\leq\sum_{Q\in{\mathcal{S}}}|Q|\left(\frac{1}{|Q|}\int_{Q}|f(x)|dx\right)\left(\frac{1}{|Q|^{t}}\int_{Q}\mbox{1\hskip-4.25pt{1}}_{E^{\prime}}(x)dx\right)$
@@
Indeed, this follows directly by Step I in Theorem 5.7 and the pointwise sparse bound $M_{t}^{{\mathcal{S}}}f(x)\lesssim\mathbb{A}_{{\mathcal{S}}}^{t}f(x),\;x\in\mathbb{R}^{n}$. Therefore, by (5.3.1), Lorentz–Hölider’s inequality, and the $L^{p,1}(\mathbb{R}^{n})$ boundedness [^7] This follows from the off-diagonal Marcinkiewicz interpolation theorem applied to the $M_{\mathrm{HL}}^{\mathcal{D}}$ (see, e.g., [17, Theorem 1.4.19]). of $M_{\textrm{HL}}^{{\mathcal{D}}}$, we have

### v6

@@ Abstract.
Motivated by the work of Cheng–Fang–Wang–Yu on the hypersingular Bergman projection, we develop a real-variable framework for hypersingular operators in regimes where strong-type bounds fail on the critical line. Our main new ingredient is the Forelli–Rudin method: a dyadic mechanism, inspired by complex–analytic Forelli–Rudin type arguments, that yields sharp critical-line and endpoint estimates.
On the unit disc, for $1<t<3/2$, we give a complete $(p,q)$-mapping characterization for the dyadic hypersingular maximal operator $\mathcal{M}_{t}^{\mathcal{D}}$, including sharp bounds on the critical line 1/q1/p=2t21/q-1/p=2t-2 and a weighted endpoint criterion in the radial setting. We also prove a novel two-weight estimate for $\mathcal{M}_{t}^{\mathcal{D}}$ in the range $p>q$, valid for all $t>0$. For the hypersingular Bergman projection
@@
we establish sharp critical-line bounds, with emphasis on the endpoint weak-type estimate at $(p,q)=\bigl(\tfrac{1}{3-2t},1\bigr)$. In particular, this result resolves an open question on the critical-line behavior of the Bergman projection in the hypersingular regime. Finally, we introduce a class of hypersingular cousins of sparse operators in $\mathbb{R}^{n}$ associated with graded sparse families, quantified by the sparseness $\eta$ and a new structural parameter (the degree) $K_{\mathcal{S}}$. We characterize the corresponding sharp strong- and weak-type regimes in terms of $(n,t,\eta,K_{\mathcal{S}})$.
This real-variable perspective addresses an inquiry of Cheng–Fang–Wang–Yu on developing effective real-analytic tools in the hypersingular regime for both ${\mathcal{M}}_{t}^{{\mathcal{D}}}$ and $K_{2t}$, and it also provides a new route to critical-line analysis for Forelli–Rudin type and related hypersingular operators in both real and complex settings.
@@ 3. $\bullet$
We have the following full characterization for ${\mathcal{M}}_{t}^{{\mathcal{D}}}$, including the critical line estimates. Let $1<t<3/2$ and $\mathcal{D}$} be any dyadic system on $\mathbb{T}$. Then ${\mathcal{M}}_{t}^{{\mathcal{D}}}$ (as well as ${\mathcal{M}}_{t}$) is
@@
Moreover, the above estimates are sharp (see, Proposition 3.9). The results above are summarized in Figure 1 below.
@@
Moreover, the above estimates for $K_{2t}$ are sharp (see, Proposition 4.2). We summarize the above results for $K_{2t}$ in Figure 2 below.
@@
${\mathcal{S}}_{m}:=\left\{I_{k}:=\left[\frac{k}{2^{m}},\frac{k+1}{2^{m}}\right),0\leq k\leq 2^{m}-1\right\}\cup\left\{[0,2)\right\},$
@@
which implies $\left\|\mathbb{A}_{{\mathcal{S}}_{m}}^{t}\right\|_{L^{\infty}([0.2))\to L^{1}([0,2))}\gtrsim 2^{m(t-1)}$. This example shows that, in order to capture the behavior of $\mathbb{A}_{{\mathcal{S}}}^{t}$, the sparseness of $\mathcal{S}$} alone does not suffice, and one must have certain control for the dyadic scales between consecutive layers in $\mathcal{S}$}. This simple example also highlights a fundamental difference between the sparse operator and its hypersingular counterpart: in the former case, the degree plays essentially no role. We refer the reader to Section 5.1 for further discussion of this topic.
@@
The difficulties at this endpoint are two–fold:
@@
This identity serves as a bridge between analysis on holomorphic function spaces and dyadic real-variable harmonic analysis. Indeed, its left-hand side is a boundary distance from the intrinsic geometry of $\mathbb{D}$}, while its right-hand side is a counting quantity for dyadic Carleson boxes. It allows us to replace the boundary scale by a dyadic level, and it is the starting point for the dyadic Forelli–Rudin method.
@@
    The above estimates for $\mathbb{A}_{{\mathcal{S}}}^{t}$ are sharp, as is already reflected in our earlier strong and weak-type characterizations of ${\mathcal{M}}_{t}^{{\mathcal{D}}}$ and $K_{2t}$. These observations suggest that the $L^{p}$$L^{q}$ mapping properties of $K_{2t}$ are governed primarily by the geometry of (dyadic) Carleson boxes, rather than analyticity.
@@
    From the viewpoint of complex function theory, we develop a systematic theory of the hypersingular maximal operator ${\mathcal{M}}_{t}^{{\mathcal{D}}}$ and also advance the study of Forelli–Rudin type operators in the hypersingular regime by establishing new sharp endpoint and critical-line estimates for $K_{2t}$ with $t>1$. To the best of our knowledge, these estimates are new even in the unit disc setting. A key feature of our approach is the introduction of Forelli–Rudin method into this setting, which allows us to overcome the lack of strong-type bounds and provides a robust harmonic-analytic framework for hypersingular operators in both real and complex settings.
@@
The rest of this paper is organized as follows. In Section 2, we collect basic definitions and notation that will be used throughout the paper. In Section 3, we establish the sharp $L^{p}$ mapping theory for the dyadic hypersingular maximal operator ${\mathcal{M}}_{t}^{\mathcal{D}}$, including both the off-critical-line bounds and the critical-line estimates, together with the endpoint characterizations in the radial weighted setting. Moreover, we prove a two-weight estimate for ${\mathcal{M}}_{t}^{{\mathcal{D}}},t>0$. In Section 4, we turn to the hypersingular Bergman projection $K_{2t}$ and prove the sharp critical-line estimates, with special emphasis on the endpoint $(p,q)=\left(\frac{1}{3-2t},1\right)$ via a complex–analytic version of the Forelli–Rudin method. In Section 5, we introduce hypersingular sparse operators associated with graded sparse families in $\mathbb{R}^{n}$ and prove the corresponding $L^{p}$-$L^{q}$ theory via a dyadic version of the Forelli–Rudin method, which in particular extends and strengthens the results obtained earlier for ${\mathcal{M}}_{t}^{{\mathcal{D}}}$ and $K_{2t}$. In addition, we include a discussion of two other different approaches to study $\mathbb{A}_{{\mathcal{S}}}^{t}$: one via ideas from sparse domination, and the other via Bourgain’s interpolation trick. Finally, in Section 6, we propose several open problems and further directions motivated by the methods developed in this paper.
@@ 2. Notations
$\left\|f\right\|_{L^{p,r}({\mathbb{D}})}:=\left(p\int_{0}^{\infty}\lambda^{r}\left|\left\{z\in{\mathbb{D}}:|f(z)|>\lambda\right\}\right|^{\frac{r}{p}}\frac{d\lambda}{\lambda}\right)^{\frac{1}{r}},\qquad 0<r<+\infty,$
@@ 3.2. Critical line estimate
$2^{-k}$ $\simeq\left|\bigcup_{I\in{\mathcal{D}}_{k}}Q_{I}\right|\lesssim\left|\left\{w\in{\mathbb{D}}:{\mathcal{M}}_{t}^{{\mathcal{D}}}f_{k}(w)>\widetilde{c}2^{k(2t-1)}\cdot\int_{1-\frac{1}{2^{k}}}^{1-\frac{1}{2^{k+1}}}\frac{dr}{\omega^{\frac{3-2t}{2t-2}}(r)}\right\}\right|$
@@ 2. $\bullet$
To this end, we show that the $L^{p}$ estimates we derived for ${\mathcal{M}}_{t}^{{\mathcal{D}}}$ are sharp.
Proposition 3.9.
For any $1\leq p,q\leq+\infty$ with $\frac{1}{q}-\frac{1}{p}<2t-2$, ${\mathcal{M}}_{t}^{{\mathcal{D}}}$ maps $L^{p}({\mathbb{D}})$ unboundedly to $L^{q,\infty}({\mathbb{D}})$.
Proof.
To begin with, we may assume that $p,q<+\infty$. The case $p=+\infty$ follows by an argument similar to that in Observation 3.1, while the case $q=+\infty$ is trivial.
Take $\varepsilon>0$ sufficiently small, such that $1/q-1/(p+\varepsilon)<2t-2$, and let the test function $f(z):=(1-|z|^{2})^{-\frac{1}{p+\varepsilon}},\;z\in{\mathbb{D}}$. On one hand,
$\left\|f\right\|_{L^{p}({\mathbb{D}})}^{p}\simeq\int_{0}^{1}\frac{dr}{(1-r)^{\frac{p}{p+\varepsilon}}}<+\infty.$
On the other hand,
${\mathcal{M}}_{t}^{{\mathcal{D}}}f(z)$ $=\sup_{z\in Q_{I},\;I\in{\mathcal{D}}}\frac{1}{|Q_{I}|^{t}}\int_{Q_{I}}\frac{1}{(1-|w|^{2})^{\frac{1}{p+\varepsilon}}}dA(w)$
$\geq\sup_{z\in Q_{I},\;I\in{\mathcal{D}}}\frac{1}{|Q_{I}|^{t}}\int_{Q_{I}^{\textrm{up}}}\frac{1}{(1-|w|^{2})^{\frac{1}{p+\varepsilon}}}dA(w)$
$\simeq\sup_{z\in Q_{I},\;I\in{\mathcal{D}}}\frac{1}{|Q_{I}|^{t-1+\frac{1}{2(p+\varepsilon)}}}\simeq\frac{1}{(1-|z|^{2})^{2(t-1)+\frac{1}{p+\varepsilon}}},$
which gives
$\left|\left\{z\in{\mathbb{D}}:{\mathcal{M}}_{t}^{{\mathcal{D}}}f(z)\gtrsim\lambda\right\}\right|\gtrsim\min\left\{1,\lambda^{-\frac{1}{q\left(2t-2+\frac{1}{p+\varepsilon}\right)}}\right\}.$
Hence,
$\left\|{\mathcal{M}}_{t}^{{\mathcal{D}}}f\right\|_{L^{q,\infty}({\mathbb{D}})}=\sup_{\lambda>0}\lambda\left|\left\{z\in{\mathbb{D}}:{\mathcal{M}}_{t}^{{\mathcal{D}}}f(z)\gtrsim\lambda\right\}\right|^{\frac{1}{q}}\gtrsim\sup_{\lambda>0}\lambda\cdot\min\left\{1,\lambda^{-\frac{1}{q\left(2t-2+\frac{1}{p+\varepsilon}\right)}}\right\}=+\infty,$
where in the last estimate above, we used the fact that $q\left(2t-2+\frac{1}{p+\varepsilon}\right)>1$. ∎
@@ 3.3. A two-weight estimate for $\mathcal{M}_{t}^{\mathcal{D}}$
Let $1\leq q<p\leq+\infty$ and $t>0$. Let $\mathcal{D}$ be a dyadic system on $\mathbb{T}$. Let $\mu$ and $\omega$ be two weights on $\mathbb{D}$, and set $\sigma:=\omega^{-1/(p-1)}$. Assume that $\mu,\sigma\in\mathbf{B}_{\infty}({\mathcal{D}})$. Then the following statements are equivalent.
@@
$(1)\Longrightarrow(2)$.   Let $\{a_{I}\}_{I\in{\mathcal{D}}}$ be any sequence of non-negative numbers, and
@@
${\mathcal{M}}_{t}^{{\mathcal{D}}}F(z)\geq\frac{1}{|Q_{I}|^{t}}\int_{Q_{I}}|F(z)|dA(z)\geq\frac{1}{|Q_{I}|^{t}}\int_{Q^{\textrm{up}}_{I}}|F(z)|dA(z)=\frac{a_{I}\sigma(Q_{I}^{\textrm{up}})}{|Q_{I}|^{t}}\gtrsim\frac{a_{I}\sigma(Q_{I})}{|Q_{I}|^{t}},$
@@
$\int_{{\mathbb{D}}}\left|{\mathcal{M}}_{t}^{{\mathcal{D}}}F(z)\right|^{q}\mu(z)dA(z)$ $\gtrsim\sum_{I\in{\mathcal{D}}}\int_{Q_{I}^{\textrm{up}}}\left(\frac{a_{I}\sigma(Q_{I})}{|Q_{I}|^{t}}\right)^{q}\mu(z)dA(z)$
@@ 4. Critical line estimates for the hypersingular Bergman projection
(4.1) $\left\{\left(\frac{1}{p},\frac{1}{q}\right)\in[0,1]^{2}:\frac{1}{q}-\frac{1}{p}=2t-2\right\}.$
@@
As a consequence, we show that $K_{2t}$ satisfies a weak-type bound only in the region on and above the critical line (4.1).
Proposition 4.2.
For any $1<t<3/2$ and $1\leq p,q\leq+\infty$ with $\frac{1}{q}-\frac{1}{p}<2t-2$, $K_{2t}$ maps $L^{p}({\mathbb{D}})$ unboundedly to $L^{q,\infty}({\mathbb{D}})$.
Proof.
Without loss of generality, we may assume that $q<+\infty$. The unboundedness in the case $q=+\infty$ follows immediately from [14, Theorem 3].
Assume that $K_{2t}$ maps $L^{p}({\mathbb{D}})$ boundedly into $L^{q,\infty}({\mathbb{D}})$. First, we consider the case $q>1$. Fix $\varepsilon>0$ sufficiently small so that $q-\varepsilon>1$ and $1/(q-\varepsilon)-1/p<2t-2$. Then using the fact that $L^{q,\infty}({\mathbb{D}})\subseteq L^{q-\varepsilon}({\mathbb{D}})$ (see, e.g., [17, Exercise 1.1.11]), we have
$\|K_{2t}f\|_{L^{q-\varepsilon}({\mathbb{D}})}\lesssim\|K_{2t}f\|_{L^{q,\infty}({\mathbb{D}})}\lesssim\|f\|_{L^{p}({\mathbb{D}})}.$
This contradicts [14, Theorem 3].
[…]
@@
Therefore, it remains to verify claims (1) and (2). Claim (1) follows from [35, Theorem 3] by choosing $a=2t-2,b=0,c=2t,n=1,\alpha=\beta=0$, and p=q=1/(32t)p=q=1/(3-2t) there. Claim (2) follows from a direct computation. ∎
@@ 5. Hypersingular sparse operators
    $E(Q)\cap Q^{\prime}=\emptyset$ for any $Q^{\prime}\subsetneq Q$, $Q^{\prime}\in{\mathcal{S}}$.
@@ 5.1. Graded family
We now turn to some details. Let ${\mathcal{G}}\subseteq{\mathcal{D}}$ be a collection of dyadic cubes. Again, we may assume $Q\subseteq[0,1]^{n}$ for all $Q\in{\mathcal{G}}$.
@@
We call $K_{{\mathcal{G}}}$ the degree of $\mathcal{G}$}, and we refer to ${\mathcal{G}}^{(j)}$ as the $j$-th layer of $\mathcal{G}$}.
@@
    Here, we may assume that ${\mathcal{G}}^{(1)}\neq\emptyset$. Otherwise, $\mathbb{A}_{{\mathcal{G}}}^{t}f(x)=\mbox{1\hskip-4.25pt{1}}_{[0,1]^{n}}(x)\int_{[0,1]^{n}}f$, which is a rank-one operator and, in particular, maps $L^{1}(\mathbb{R}^{n})$ boundedly into $L^{\infty}(\mathbb{R}^{n})$.
@@
    Here are some examples of graded family. Consider the collection of all dyadic cubes contained in $0,1]^{n}$ together; then such a collection of dyadic cubes is graded with degree $1$. Another example is the collection [^5] In that setting, the role of $Q_{0}$ is replaced by $\mathbb{D}$}, and the sidelength $\ell(Q)$ in Definition 5.5 is replaced by the length of the boundary arc associated with the Carleson box. of dyadic Carleson boxes that appeared in our earlier analysis of ${\mathcal{M}}_{t}^{{\mathcal{D}}}$ and $K_{2t}$. This family is also graded, again with degree $1$. Finally, we observe that $K_{{\mathcal{G}}}\geq 1$.
@@
    (Critical line estimate) $\mathbb{A}_{{\mathcal{S}}}^{t}$ extends to a bounded operator from $L^{p}(\mathbb{R}^{n})$ to $L^{q,\infty}(\mathbb{R}^{n})$ when $(p,q)$ belongs to
@@
Theorem 5.7 strengthens our earlier results on the boundedness behavior of $K_{2t}$ (see, Section 4). Moreover, the proof of Step II in Theorem 5.7 yields an alternative proof of [14, Theorem 3], and the proof of Step III in in Theorem 5.7 offers a different proof of Proposition 4.3. Moreover, Theorem 5.7 is sharp, since it recovers the $L^{p}$ theory for ${\mathcal{M}}_{t}^{{\mathcal{D}}}$ and $K_{2t}$ as special cases.
@@
respectively. Note that the set $\{x\in[0,1]^{n}:{\bf N}(x)=+\infty\}$ has measure zero. We have the following claims.
@@
$\mathbb{A}_{{\mathcal{S}}}^{t}f(x)\lesssim{\bf W}(x)M_{\textrm{HL}}^{{\mathcal{D}}}f(x),\qquad a.e.\quad x\in[0,1]^{n},$
@@
(5.9) $\mathbb{A}_{{\mathcal{S}}}^{t}f(x)=\sum_{x\in Q\in{\mathcal{S}}}|Q|^{1-t}\cdot\frac{1}{|Q|}\int_{Q}|f(y)|dy\leq M_{\textrm{HL}}^{{\mathcal{D}}}f(x)\cdot\sum_{x\in Q\in{\mathcal{S}}}|Q|^{1-t},\qquad a.e.\quad x\in[0,1]^{n}.$
@@
(5.10) $\sum_{x\in Q\in{\mathcal{S}}}|Q|^{1-t}\leq\sum_{j=0}^{{\bf N}(x)}2^{njK_{{\mathcal{S}}}(t-1)}\lesssim 2^{n{\bf N}(x)K_{{\mathcal{S}}}(t-1)}={\bf W}(x),\qquad a.e.\quad x\in[0,1]^{n}.$
@@
and let $f\in L^{p,1}(\mathbb{R}^{n})$. Without loss of generality, we may assume $\textrm{supp}f\subseteq[0,1]^{n}$. By [17, Exercise 1.4.14] and using the assumption that ${\mathbb{A}}_{{\mathcal{S}}}^{t}$ is localized in $0,1]^{n}$, we have
@@
$=\sum_{Q\in{\mathcal{S}}}|Q|\left(\frac{1}{|Q|}\int_{Q}|f(x)|dx\right)\left(\frac{1}{|Q|^{t}}\int_{Q}\mbox{1\hskip-4.25pt{1}}_{E^{\prime}}(x)dx\right)$
@@
Indeed, this follows directly by Step I in Theorem 5.7 and the pointwise sparse bound $M_{t}^{{\mathcal{S}}}f(x)\lesssim\mathbb{A}_{{\mathcal{S}}}^{t}f(x),\;x\in\mathbb{R}^{n}$. Therefore, by (5.3.1), Lorentz–Hölider’s inequality, and the $L^{p,1}(\mathbb{R}^{n})$ boundedness [^7] This is a standard consequence of the off-diagonal Marcinkiewicz interpolation theorem applied to $M_{\mathrm{HL}}^{\mathcal{D}}$ (see, e.g., [17, Theorem 1.4.19]). of $M_{\textrm{HL}}^{{\mathcal{D}}}$, we have